OpenStudy (anonymous):

What is the slope of the line that passes through the points (–6,1) and (4,–4)? A. -2 B. 2 C. -½ D. ½

OpenStudy (nubeer):

m = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what do i do next

OpenStudy (amistre64):

well, if youve used the formula correctly, you choose the option that matches it

OpenStudy (nubeer):

plug in the values u have of x1,x2,y1,y2

OpenStudy (amistre64):

personally, i step thru it subtract one point from the other and stack y over x

OpenStudy (amistre64):

(–6, 1) -( 4,–4) ------- x , y ; slope = y/x

OpenStudy (anonymous):

do i subtract

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

(–6, 1) -( 4,–4) ------- x , y what do you get for x and y?

OpenStudy (anonymous):

2=x y=1

OpenStudy (amistre64):

(–6, 1) -( 4,–4) ------- -10, 5

OpenStudy (anonymous):

ok i c

OpenStudy (amistre64):

now just stack y/x and simplify the fraction

OpenStudy (anonymous):

negative one half

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

thank you for yur help

OpenStudy (amistre64):

i have to do it in that way, becasue i tend to put the wrong values into the wrong places inthe formula. subtract and stack is much simpler for me to keep things straight :) good luck

OpenStudy (anonymous):

thank you